OpenStudy (anonymous):

Determine Laplace transform of f(t)=10sin(2t-30 degrees) ?

OpenStudy (anonymous):

I am unsure about the degree shift, but I know that \[L(\sin(2t))=2/(s^2+4)\] And so \[L(10\sin(2t))=20/(s^2+4)\] I hope this helped

OpenStudy (anonymous):

you could do the following first: sin(2t-30)=sin2t cos30 + cos2t sin30=... after that - Laplace transform...

OpenStudy (anonymous):

check this one - big help!!

OpenStudy (anonymous):

Thank you for -pdf.